how do i solve for f(x)=81-x^2
Answer by Theo(13342)

set f(x) = 0 and factor it as follows:
start with:
81 - x^2 = 0
add x^2 to both sides of the equation to get:
81 = x^2
take the square root of both sides to get:
x = +/- sqrt(81) which results in:
x = +/- 9
9^2 = 81
(-9)^2 = 81